# OutLines: Modeling Spectral Lines from Winds, Bubbles, and Outflows

### v1

@@ A.3 Cavity Effects
Figure 22: Same as Figure 21 but for a directed outflow with an inner cavity.
@@ A.4 Obstruction by a Disk
In §2.4, we present the possible inclusion of a disk which encircles the source of the outflow and is inclined orthogonally to the cone axis. This disk obstructs the posterior cone but may affect either the red-shifted or both wings of the line profile. OutLines projects the disk onto the line of sight and then determines the length of each deprojected velocity which is still observed after accounting for the disk effect. Since the disk is centered at the source of the outflow and assumed to be orthogonal to the outflow cone, the projected disk is simply an ellipse with semi-major and semi-minor axes of ($w[x_{d}|\beta],w[x_{d}|\beta]\sin i$). Thus, the intersection points ($p_{d}$$q_{d}$) between the deprojected velocity and the disk can be written as
$p_{d}=\pm\sqrt{r_{\rm b}^{2}-q_{d}^{2}}$ (51)
@@
$q_{d}=\pm\sqrt{\frac{r_{\rm b}^{2}-w^{2}[x_{d}|\beta]}{\csc^{2}i-1}}.$ (52)
Rather than compute $\ell_{\rm w}$ and $\ell_{\rm disk}$ separately, which would needlessly incur a computational overhead, we can instead simply compare $q$ and $q_{d}$ to determine the final $\ell_{\rm w}$. For the posterior cone, $\ell_{\rm w}$ is given by
$\pi\ell_{\rm w}=\begin{cases}\arccos{\frac{q}{r_{\rm b}}}&q>q_{d}\\ \arccos{\frac{q_{d}}{r_{\rm b}}}&-q_{d}<q<q_{d}\\ 2\arccos{\frac{q_{d}}{r_{\rm b}}}-\arccos{\frac{q}{r_{\rm b}}}&q<-q_{d}\\ \end{cases}.$ (53)
@@
Appendix B Non-zero Launch Velocities
$w_{\rm nfr}=w[x|\beta]\left(1-\frac{v_{0}}{v_{\infty}}\right)+\frac{v_{0}}{v_{\infty}}.$ (55)
Emergent profiles are similar to those presented in TenorioTagle1996, which included distributions of non-zero shell ages, and to those presented in 2015ApJ...801...43S, which account for non-zero launch velocities. We show example line profiles with varying launch velocities in Figure 23. Future work will revisit this possibility in greater detail.
Figure 23: Effect of varying the launch velocity from rest ($v_{0}=0$) to 30% of the terminal velocity on nebular emission (left) and absorption (right) line profiles, here normalized to the case of launching from rest for visualization.
Table 2: Left: the Profile object classes in OutLines with which users can generate model profiles for different spectral line types. Right: keyword settings for the Profile classes to customize models.
@@ C.1 Installation
C.2 The Profile Classes
The primary front-end user interaction with OutLines is through the various Profile object classes, which we list in Table 2. Calling one of these classes instantiates an object containing the line profile model, parameter guesses, and parameter bounds. When calling a class, users need only provide the relevant atomic data and any optional keyword arguments specifying the velocity field, geometry, and density profile to customize the model (see Table 2). The terms and related keyword arguments for the profile model are listed in Table 3 for the geometry, Table 4 for the velocity field, and Table 5 for the density profile, with the OutLines defaults indicated by a .
Table 3: Geometry keyword options and related required arguments for OutLines. indicates the OutLines default. indicates geometry only applicable to emission lines.
@@ C.2 The Profile Classes
‘BetaCAK’  15 1975ApJ...195..157C; Castor1979
‘AccPlaw’ 16 2010ApJ...717..289S
‘VelPlaw’ 17 FaucherGiguere2012
@@
‘PowerLaw’  $\alpha$ 18
‘Exponential’ $\gamma$ 19
‘Normal’ $x_{1}$, $\sigma_{x}$ 21
‘LogNormal’ $\log x_{1}$, $\log\sigma_{x}$ 22
‘Shell’ $x_{1}$, $\sigma_{x}$ 24
‘FRED’ $x_{1}$, $\tau_{1}$, $\tau_{2}$ 23
‘DoublePowerLaw’ $\alpha_{1}$, $\alpha_{2}$, $x_{1}$ 20
‘Pulses’ $x_{1}$, $\sigma_{x}$, $x_{k}$ 25
‘DampedPulses’ $\gamma$, $x_{1}$, $\sigma_{x}$, $x_{k}$ 26
Users will primarily use OutLines by instantiating one of the several Profile classes: Nebular, Resonant, Fluorescent, and Absorption. The object-bound methods of the Profile classes can be used to obtain or update parameter values and to generate line profiles. We provide examples of using these classes to generate line profiles in Appendix D. To obtain line profiles computed by OutLines, users can call the get_profile() bound method for either class, which requires only an array of wavelengths and computes a line profile for the parameters stored in the object. Alternatively, the line profile function is directly accessible through the Profile attribute, which requires a wavelength array just as get_profile() but also takes the full set of parameters. For purposes of modeling observations, users may prefer to use the Profile function where the input parameters may be readily varied without updating the corresponding attribute in the object.
@@ C.3 The Properties Class
v.out 32 $v_{out}$ characteristic radius
Rcal 33 $\mathcal{R}$ relative column density
Mdot 37 $\dot{M}/n_{0}R_{0}^{2}$ mass outflow rate
pdot 38 $\dot{p}/n_{0}R_{0}^{2}$ momentum flux
Edot 39 $\dot{E}/n_{0}R_{0}^{2}$ mechanical luminosity
C.4 Visualization
@@
To assist with interpretation of outflow properties and geometry, OutLines contains several functions to generate figures like the ones presented here. To produce isocontours as in Figure 6, we provide the PlotIsoContours function. To produce the edge-on geometric cartoon as in Figure 10, we provide the PlotGeometry function. To produce the velocity projections for filled and hollow cones as in Figures 21 and 22, we provide the PlotConeProjection function.
C.5 Numerical Methods
@@
Some equations require root-finding where analytic solutions are not possible, including the “source-effect” and aperture velocity limits (Equations 28 and 30, respectively) and the characteristic outflow radius $x_{out}$ where the momentum density is maximized. As in Flury2023, we employ Brent’s method implemented by scipy.optimize to solve for the root in each case.
@@
The initial profile model of Flury2023 solved the integral for the emission line profile (Equation 10) using the adaptive Romberg method, which included Richardson extrapolation under the Bulirsch-Stoer method. As implemented by the now-deprecated scipy.integrate.romberg, this method can be costly and in later scipy versions has loss of accuracy in non-spherical geometries due to poor sampling of the integrand. For improvements to both speed and accuracy, we now adopt Gaussian quadrature with a fixed number of nodes (as implemented by scipy.integrate.fixed_quad) to solve each of the integrals for $\phi_{\lambda}$. We find at least 16 nodes necessary to obtain reliable profile solutions in many cases but conservatively assume 256 nodes for optimal accuracy across the profile parameter spaces.
@@
For a single core on a conventional laptop computer, benchmarking indicates that OutLines can calculate an emission line profile at $R=5,000$ in $\approx$20 ms for any assumed geometry. Given reasonable guess parameters, iterative non-linear least squares fitting routines like the trust-reflective algorithm (see scipy.optimize.least_squares) perform rapidly, typically converging on a solution within 10 to 30 s depending on the data and the assumed geometry. For 25 walkers, a typical MCMC sampler (here benchmarked using emcee, emcee) takes anywhere from $0.1$ to $2$ s per step to run on a conventional laptop computer when fitting an OutLines model to an observed line profile, even when accounting for complex geometries and LSF effects. The addition of convolution with a line spread function does increase the computation time; however, the extent to which this occurs depends on the resolution, size of the data set, and the implementation of the convolution.
Appendix D Example Line Profiles with OutLines
@@
D.1 C ii $\lambda 1334$
Example instantiations of Absorption, Resonant, and Fluorescent objects from the Profile class using atomic data from Morton2003. Here, the object classes are instantiated, documentation, setting, and parameters are printed to the terminal, and the column density and terminal velocity are updated.
@@ D.1 C ii $\lambda 1334$
D.2 [O iii] $\lambda 4959,5007$
Example insantiation of the Nebular object from the Profile class in OutLines using atomic data from NIST (NIST_ASD) and Storey2000. Here, we illustrate how to customize the model with keyword arguments passed to the object class on instantiation, updating parameters, printing setting and parameters, computing and printing outflow properties, and visualizing a line profile as in Figure 24.
@@
Figure 24: Example [O iii] $\lambda\lambda 4959,5007$ line doublet profile produced using OutLines as shown in Appendix D.2.
Appendix E Example Modeling
@@
E.1 Si ii $\lambda 1190,1193$ Absorption Doublet
@@
E.2 [S ii] $\lambda 6716,6731$ Emission Doublet
@@
E.3 [O iii] $\lambda 4959,5007$ Emission Doublet for the Seyfert 2 in NGC 2992
@@
7# MC sampler for posterior
8import emcee
9# log prior -- uniform bounded
10def log_prior(self,theta):
@@
15 # suggest that the best-guess angles
16 # are accurate within 10 degrees
@@
18 return -0.5*((2.99e5*theta[1]-1100)/300)**2 \
19 -0.5*((2.99e5*theta[0]- 75)/50)**2 \
20 -0.5*((theta[7]-1.25)/0.2)**2 \
21 -0.5*((theta[8]-0.95)/0.2)**2 \
22 -0.5*((theta[9]-0.40)/0.2)**2 \
23 -0.5*((theta[4]/theta[3]-2.98)/0.05)**2 \
@@
27# instantiate the nebular line
28# profile object using
29# rest wavelengths from NIST
30kwargs = dict(Geometry='HollowCones',AddStatic=True,\ VelocityField='AccPlaw',Disk=True)
31model = OL.Nebular([4958.911,5006.843],**kwargs)
32# update some parameters
33model.update_params(['TerminalVelocity'],[500,30,15,45])
34model.update_params(['OpeningAngle','CavityAngle',\ 'Inclination'],[60,30,60])
35model.update_params(['FluxOutflow1','FluxStatic1'],\ [1/2.98,1/2.98]) # Storey & Zeppen 2000
36# simple trust-reflective fit for initial guess
37tr_kwargs = dict(p0=model.get_params(),\ bounds=model.get_bounds(),sigma=flux_error)
38p0,c0 = curve_fit(model.Profile,wave,flux,**tr_kwargs)
39model.update_params(model.get_param_names(),p0)
40# initialize walkers and sampler and run the sampler
41init,nwalk = model.init_params()
42# use the custom prior
43model.log_prior = MethodType(log_prior,model)
44# run the MCMC sampler
45sampler = emcee.EnsembleSampler(nwalk,model.npar,\ model.log_probability,args=(wave,flux,flux_error))
46sampler.run_mcmc(init,1200, progress=True)
47samples = sampler.get_chain(discard=200,flat=True)
48# extract parameters from posterior sample
49for i,param in enumerate(model.get_param_names()):
@@
52# print line profile parameters to terminal
53model.print_params()
54# obtain outflow properties
55props = OL.Properties(model)
56# print outflow properties to terminal
57props.print_props()
Appendix F Posterior Sampling
All posterior samples for each of the OutLines modeling cases were obtained using emcee (emcee) and visualized with figures produced by corner (corner) and subsequently adapted. Results are shown for NGC 5471 Knot A in Figure 25, the super star cluster of Green Pea J1044+0353 in Figure 26, Mrk 1486 in 27, and the Seyfert 2 nucleus of NGC 2992 in 28.

### v2

@@ A.3 Cavity Effects
Figure 24: Same as Figure 23 but for directed cones with an inner cavity.
@@ A.4 Obstruction by a Disk
In §2.4, we present the possible inclusion of a disk which encircles the source of the outflow and is inclined orthogonally to a cone or hemisphere axis. This disk obstructs the posterior cone or hemisphere but may affect either the red-shifted or both wings of the line profile. OutLines projects the disk onto the line of sight and then determines the length of each deprojected velocity which is still observed after accounting for the disk effect. Since the disk is centered at the source of the outflow and assumed to be orthogonal to the outflow cone, the projected disk is simply an ellipse with semi-major and semi-minor axes of ($w[x_{d}|\beta],w[x_{d}|\beta]\cos i$). Thus, the intersection points ($p_{d}$$q_{d}$) between the deprojected velocity and the disk can be written as
$p_{d}=\pm\sqrt{r_{\rm w}^{2}-q_{d}^{2}}$ (55)
@@
$q_{d}=\pm\sqrt{\frac{r_{\rm w}^{2}-w^{2}[x_{d}|\beta]}{\sec^{2}i-1}}.$ (56)
Rather than compute $\ell_{\rm w}$ and $\ell_{\rm disk}$ separately, which would needlessly incur a computational overhead, we can instead simply compare $q$ and $q_{d}$ to determine the final $\ell_{\rm w}$. For the posterior cone or hemisphere, $\ell_{\rm w}$ is given by
$\pi\ell_{\rm w}=\begin{cases}\arccos{\frac{q}{r_{\rm w}}}&q>q_{d}\\ \arccos{\frac{q_{d}}{r_{\rm w}}}&-q_{d}<q<q_{d}\\ 2\arccos{\frac{q_{d}}{r_{\rm w}}}-\arccos{\frac{q}{r_{\rm w}}}&q<-q_{d}\\ \end{cases}.$ (57)
@@
Appendix B Implementation of OutLines
Table 2: Left: the Profile object classes in OutLines with which users can generate model profiles for different spectral line types. Right: keyword settings for the Profile classes to customize models. Disk not implemented for spherical geometries. Pulse only applicable ensemble density profiles.
@@ B.1 Installation
B.2 The Profile Classes
The primary front-end user interaction with OutLines is through the various Profile object classes, which we list in Table 2. Calling one of these classes instantiates an object containing the line profile model, parameter guesses, and parameter bounds. When calling a class, users need only provide the relevant atomic data and any optional keyword arguments specifying the velocity field, geometry, and density profile to customize the model (see Table 2). The terms and related keyword arguments for the profile model are listed in Table 3 for the geometry, Table 4 for the velocity field, and Table 5 for the density profile, with the OutLines defaults indicated by a . In some cases, users may want to consider a velocity field or density profile not included in OutLines. In such cases, if a user passes a function to either keyword instead of a string, OutLines will implement the user-defined field or profile.
Table 3: Geometry keyword options and related required arguments for OutLines. indicates the OutLines default.
@@ B.2 The Profile Classes
‘BetaCAK’  15 Castor et al. (1975a); Castor and Lamers (1979)
‘AccPlaw’ 16 Steidel et al. (2010)
‘VelPlaw’ 17 Faucher-Giguère and Quataert (2012)
@@
Winds
‘PowerLaw’  $\alpha$ 19
‘Exponential’ $\gamma$ 20
‘DoublePowerLaw’ $\alpha_{1}$, $\alpha_{2}$, $x_{1}$ 21
Bubbles / Shells
‘Normal’ $x_{1}$, $\sigma_{x}$ 22
‘LogNormal’ $\log x_{1}$, $\log\sigma_{x}$ 23
‘DLogic‘ $\log x_{1}$, $k$ 25
‘Shell’ $x_{1}$, $\sigma_{x}$ 26
‘FRED’ $x_{1}$, $\tau_{1}$, $\tau_{2}$ 24
Ensembles
‘Pulses’ $x_{1}$, $\sigma_{x}$, $x_{k}$ 27
‘DampedPulses’ $\gamma$, $x_{1}$, $\sigma_{x}$, $x_{k}$ 28
Users will primarily use OutLines by instantiating one of the several Profile classes: Nebular, Resonant, Fluorescent, and Absorption. The object-bound methods of the Profile classes can be used to obtain or update parameter values and to generate line profiles. We provide examples of using these classes to generate line profiles in Appendix C. To obtain line profiles computed by OutLines, users can call the get_profile() bound method for either class, which requires only an array of wavelengths and computes a line profile for the parameters stored in the object. Alternatively, the line profile function is directly accessible through the Profile attribute, which requires a wavelength array just as get_profile() but also takes the full set of parameters. For purposes of modeling observations, users may prefer to use the Profile function where the input parameters may be readily varied without updating the corresponding attribute in the object.
@@ B.3 The Properties Class
v.out 36 $v_{out}$ characteristic radius
Rcal 37 $\mathcal{R}$ relative column density
Mdot 41 $\dot{M}/n_{0}R_{0}^{2}$ mass outflow rate
pdot 42 $\dot{p}/n_{0}R_{0}^{2}$ momentum flux
Edot 43 $\dot{E}/n_{0}R_{0}^{2}$ mechanical luminosity
B.4 Visualization
@@
To assist with interpretation of outflow properties and geometry, OutLines contains several functions to generate figures like the ones presented here. To produce isocontours as in Figure 6, we provide the PlotIsoContours function. To produce the edge-on geometric cartoon as in Figure 10, we provide the PlotGeometry function. To produce the cone and velocity projection geometry for filled cones as in Figure 22, we provide the PlotProjGeom function. To produce the velocity projections for filled and hollow cones as in Figures 23 and 24, we provide the PlotConeProj function.
B.5 Numerical Methods
@@
Some equations require root-finding where analytic solutions are not possible, including the “source-effect” and aperture velocity limits (Equations 32 and 34, respectively) and the characteristic outflow radius $x_{out}$ where the momentum density is maximized. As in Flury et al. (2023), we employ Brent’s method implemented by scipy.optimize to solve for the root in each case.
@@
The initial profile model of Flury et al. (2023) solved the integral for the emission line profile (Equation 10) using the adaptive Romberg method, which included Richardson extrapolation under the Bulirsch-Stoer method. As implemented by the now-deprecated scipy.integrate.romberg, this method can be costly and in later scipy versions has loss of accuracy in non-spherical geometries due to poor sampling of the integrand. For improvements to both speed and accuracy, we now adopt Gauss-Legendre quadrature with a fixed number of nodes to solve the integrals for $\phi_{\lambda}$. We implement quadrature integration following the scipy.integrate.fixed_quad; however, for speed, we have pre-computed the nodes and weights using scipy.special.roots_legendre, which takes $\approx 1\rm\>ms$ for each call to scipy.integrate.fixed_quad. We find at least 16 nodes necessary to obtain reliable profile solutions in many cases but conservatively assume 96 nodes for optimal accuracy across the profile parameter spaces without significant computational expense. For bubbles, we restrict velocity bounds where possible to improve sampling of the density profile. Future work may include adopting Gauss-Kronrod integration for improved accuracy and precision, e.g., following Laurie (1997).
To normalize the line profile, we use a linear integration method to effectively handle cusps which may occur at peaks and troughs, typically near the line core. We implement this integration following the shoelace algorithm case of Green’s theorem. Benchmarking against the scipy.integrate.trapezoid implementation of trapezoid rule integration, we find comparable results with a marginally faster (by $1\rm\>ms$) compute time.
@@
For a single core on a conventional laptop computer, benchmarking indicates that OutLines can calculate an emission line profile at $R=5,000$ in $6\rm\>ms$ for the most complex geometry (disk plus hollow cones), and that the compute time increases almost linearly with $R$ such that run times are $2.5\rm\>ms$ at $R=2,000$, $4\rm\>ms$ at $R=3,000$, $13\rm\>ms$ at $R=10,000$, $25\rm\>ms$ at $R=20,000$, and $65\rm\>ms$ at $R=50,000$. For the simplest geometry (spherical), the compute times are much faster ($1\rm\>ms$ for $R\leq 30,000$). Given reasonable guess parameters, iterative non-linear least squares fitting routines like the trust-reflective algorithm (see scipy.optimize.least_squares) perform rapidly, typically converging on a solution within 10 to 30 s depending on the data and the assumed model. For 25 walkers, a typical MCMC sampler (here benchmarked using emcee, Foreman-Mackey et al. 2013) takes anywhere from $0.1$ to $2$ s per step to run on a conventional laptop computer when fitting an OutLines model to an observed line profile, even when accounting for complex geometries and LSF effects. The addition of convolution with a line spread function does increase the computation time; however, the extent to which this occurs depends on the resolution, size of the data set, and the implementation of the convolution.
Appendix C Example Line Profiles with OutLines
@@
C.1 C ii $\lambda 1334$
Example instantiations of Absorption, Resonant, and Fluorescent objects from the Profile class using atomic data from Morton (2003). Here, the object classes are instantiated, documentation, setting, and parameters are printed to the terminal, and the column density and terminal velocity are updated.
@@ C.1 C ii $\lambda 1334$
C.2 [O iii] $\lambda 4959,5007$
Example insantiation of the Nebular object from the Profile class in OutLines using atomic data from NIST (Kramida et al., 2024) and Storey and Zeippen (2000). Here, we illustrate how to customize the model with keyword arguments passed to the object class on instantiation, updating parameters, printing setting and parameters, computing and printing outflow properties, and visualizing a line profile as in Figure 25.
@@
Figure 25: Example [O iii] $\lambda\lambda 4959,5007$ line doublet profile produced using OutLines as shown in Appendix C.2.
Appendix D Example Modeling
@@
D.1 Si ii $\lambda 1190,1193$ Absorption Doublet
@@
D.2 [S ii] $\lambda 6716,6731$ Emission Doublet
@@
D.3 [O iii] $\lambda 4959,5007$ Emission Doublet for the Seyfert 2 in NGC 2992
@@
7# scipy convolution function for speed
8from scipy.signal import convolve
9# MC sampler for posterior
10import emcee
11# log prior -- uniform bounded
12def log_prior(self,theta):
@@
18 return -0.5*((theta[4]/theta[3]-2.98)/0.05)**2 \
@@
22# log likelihood for chi squared
23# including convolution of line profile with LSF
24# assuming the LSF kernel `kern` is defined
25def log_likelihood(self,theta,x,y,yerr):
26 ybar = convolve(model.Profile(x,*theta)+cont[ind],kern,'same')
27 ln_prb = -((y-ybar)/yerr)**2 - 2*log(yerr)
28 return nansum(ln_prb)
29# instantiate the nebular line
30# profile object using
31# rest wavelengths from NIST
32kwargs = dict(Geometry='HollowCones',AddStatic=True,\ VelocityField='AccPlaw',Disk=True)
33model = OL.Nebular([4958.911,5006.843],**kwargs)
34# update some parameters
35model.update_params(['TerminalVelocity'],[500,30,15,45])
36model.update_params(['OpeningAngle','CavityAngle',\ 'Inclination'],[60,30,60])
37model.update_params(['FluxOutflow1','FluxStatic1'],\ [1/2.98,1/2.98]) # Storey & Zeppen 2000
38# simple trust-reflective fit for initial guess
39tr_kwargs = dict(p0=model.get_params(),\ bounds=model.get_bounds(),sigma=flux_error)
40p0,c0 = curve_fit(model.Profile,wave,flux,**tr_kwargs)
41model.update_params(model.get_param_names(),p0)
42# initialize walkers and sampler and run the sampler
43init,nwalk = model.init_params()
44# use the custom prior and likelihood
45model.log_prior = MethodType(log_prior,model)
46model.log_likelihood = MethodType(log_likelihood,model)
47# run the MCMC sampler
48sampler = emcee.EnsembleSampler(nwalk,model.npar,\ model.log_probability,args=(wave,flux,flux_error))
49sampler.run_mcmc(init,1200, progress=True)
50samples = sampler.get_chain(discard=200,flat=True)
51# extract parameters from posterior sample
52for i,param in enumerate(model.get_param_names()):
@@
55# print line profile parameters to terminal
56model.print_params()
57# obtain outflow properties
58props = OL.Properties(model)
59# print outflow properties to terminal
60props.print_props()
Appendix E Posterior Sampling
All posterior samples for each of the OutLines modeling cases were obtained using emcee (Foreman-Mackey et al., 2013) and visualized with figures produced by corner (Foreman-Mackey, 2016) and subsequently adapted. Results are shown for NGC 5471 Knot A in Figure 26, the super star cluster of Green Pea J1044+0353 in Figure 27, Mrk 1486 in 28, and the Seyfert 2 nucleus of NGC 2992 in 29.